OpenStudy (anonymous):

find the equation of the line passing through the points P(1,1)and Q (4,-1) trying to figure out how to put answer in standard form

OpenStudy (anonymous):

The slope is -2/3 first put it into point slope form with the point (1,1)

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

I got the -2/3

OpenStudy (anonymous):

but it says to put answer in standard form

OpenStudy (anonymous):

where is the 5 coming from

OpenStudy (anonymous):

so y-1 = -2/3(x-1) distribute and isolate y

OpenStudy (anonymous):

k Ive got that

OpenStudy (anonymous):

it gives me these answers to choose from a) 2x-3y=-5 b)2x+3y=5 c)3x+2y=5 d)3x-2y=-5

OpenStudy (anonymous):

y=-2/3 x+5/3 since the entire right side is divided by 3, multiply the enire equation by 3 to get closer to standard form.

OpenStudy (anonymous):

If standard form is Ax+By=C, then the normal form is usually y=-A/B x + C/B

OpenStudy (anonymous):

use that and see what you get

OpenStudy (anonymous):

still trying to figure out where the 5 is coming from?

OpenStudy (anonymous):

nevermind I will figure it out

OpenStudy (texaschic101):

Find the slope using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (1,1) x1 = 1 and y1 = 1 (4,-1) x2 = 4 and y2 = -1 now lets sub slope(m) = (-1-1) / (4 - 1) m = -2/3 Now we have the slope and a set of points so we will use the point slope formula y - y1 = m(x - x1) slope(m) = -2/3 (4,-1) x1 = 4 and y1 = -1 now lets sub y - (-1) = -2/3(x - 4) y + 1 = -2/3x + 8/3 y = -2/3x + 8/3 - 1 y = -2/3x + 8/3 - 3/3 y = -2/3x + 5/3 -- this is slope intercept form now we have to turn it from slope intercept to standard form. standard form is Ax + By = C y = -2/3x + 5/3 (In standard form there is no fractions, therefore we have to multiply the equation by the LCD, which is 3, to get rid of the fractions. 3y = -2x + 5 (add 2x to both sides) 2x + 3y = 5 --> your equation in standard form

OpenStudy (anonymous):

yep @texaschic101 that's exactly it.